OpenStudy (highschoolmom2010):

does 15,20,25 for a pythagorean triple

OpenStudy (anonymous):

Apply the Pythagorean Theorem: \[15^2 + 20^2 = 25^2\]If this statement is true, it works.

OpenStudy (highschoolmom2010):

225+400=625 so im assuming it is